# Thermodynamic & pattern-matching efficiency in heterogeneous networks

### v2

@@ 3 Methods
Our results are evaluated over subsets of three different datasets: the NIST [25], the MNIST [26] and the FashionMNIST [27] datasets. These datasets contain randomly collected instances of image samples out of 10 classes. It should be noted that in this type of data, a pattern class can generally not be represented by a single centroid vector, but will contain multiple densities (’sub-style patterns’) such that the sample point cloud can be ordered naturally in a network structure. The MNIST database contains $24\times 24$ grey-scale images of handwritten digits [0-9] commonly used to benchmark machine- and deep- learning algorithms. The NIST dataset is the ancestor of the MNIST, it contains $16\times 16$ black-and white images of the handwritten digits [0-9]. The FashionMNIST is a large dataset of $28\times 28$ gray-scale images of fashion products from 10 categories. In all our experiments a subset of 2000 samples is extracted from the train-set and the test-set independently. In the Supplementary we show also the same results for larger networks generated from $4000$ samples from the MNIST and FashionMNIST. The 10 classes are equally represented both in the test and in training set. We limit to $r\leq 0.85$ as the accuracy, and hence the pattern-matching efficiency quickly drops for larger $r$, moreover the algorithm shows instability as it doesn’t necessarily converge for $r\geq 0.9$. The Portegys algorithm we use is the extended version with $k=5$ introduced in [13] and where distance between patterns is the Euclidean distance.
@@ 3.1 Pattern-matching efficiency
After the tree formation described in Section 1, the Portegys algorithm enters the search phase, where the best match to a test pattern is given using a first-best approach. The original goal of this algorithm was to restructure a list of examples in a tree-based way, such that the matching accuracy would not be much worse than nearest-neighbor matching on a full list, while providing a considerable speed up due to the tree-based structure. The search process is an iterative process which begins with the computation of the distance between the test pattern and the root-node pattern scaled by $r$. The distance is then compared to the distance between the test pattern and all of the root node’s children. The first node with a pattern that has a smaller distance is than selected as the new root node, and the process starts over with the newly selected node as root. The search phase terminates when no node whose pattern has smaller distance to the test pattern than the current one can be found.
It is natural to define an pattern-matching efficiency, $\theta$, in terms of class match accuracy, $A$, and the number of distances computed, $N_{dist}$:
@@
the logarithm is taken since the range of $N_{dist}$ can virtually span the interval $N_{dist}\in(0,N]$ where $N$ is the number of nodes. For the case of a star graph, $r\rightarrow 0$, the number of computed distances is maximum $N_{dist}=N$, for lower values of $r$ the number of computed distances decreases as does the accuracy but at a different pace, thus, it is possible to identify an optimal trade-off between the gain in computational burden and the loss in performance, as depicted in panel (c) of Figure 3.
@@
Figure 3: Pattern-matching efficiency, (c), is defined as the trade-off between the classification accuracy, (a), and the logarithm of the number of distances computed to find the best match, (b). The dashed-red vertical line indicates the optimal $r$ value in terms of efficiency. The shaded grey area designates the standard deviation. The purple shaded area in panel (c) indicates a tolerance $r$-domain of optimal efficiency. The tolerance domain is defined as the $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the maximum $\theta$, whose y-axis projection is indicated by yellow shaded area. Curves are computed over 50 independent simulations with a subset of 2000 samples extracted from the NIST dataset. The 10 digit-classes are equally represented both in the train- and test-set.
@@ 4 Results
4.1 Thermodynamic and pattern-matching efficiency at the emergence of new C-peaks
@@
Figure 4 shows the coherence between algorithmic defined in Section 3 and thermodynamic efficiency in Equation 5 for the three investigated datasets. In panels (a)-(c) the pattern-matching efficiency, $\theta$, is depicted. The vertical red-dashed line indicates the $r$ value for the optimal algorithmic trade-off, while the shaded purple region indicates a $r$-domain tolerance of the optimal pattern-matching efficiency. The tolerance domain is defined as the $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the best trade-off. In panels (d)-(f) the thermodynamic efficiency, $\eta$, shows remarkable overlap with the pattern-matching efficiency $\theta$ in terms of optimal $r$. See the $r$-range spanned by the purple-shaded areas in panels (a)-(c) and the vertical dashed lines in panels (d)-(f). The overlap between the two efficiencies suggests that, whether thermodynamic or algorithmic, the two objects are the same thing just viewed from different angles. See the next Section for the discussion on the double peak of the $\eta$ curves for the MNIST.
@@
Similar to what would be the transition from vapor/liquid to an amorphous solid, at a critical value of $r$ topological complexity increases from the trivial traits of a star graph. Thus, the emergence of structure is reflected by the arise of a second peak in $C$ together with a roughly $C=c_{0}>0$ plateau in between the first and the last peak as is illustrated in panels (e)-(i). Such transition is also depicted in Figure 5 by the bifurcation of the blue and green curves for $\tau^{(1)}_{peak}$ and for $\tau^{(2)}_{peak}$ indicating the lower and upper branches in panels (g)-(i) of Figure 4.
Before the bifurcation the two curves overlap and the single peak in the specific heat identifies a single trivial resolution scale below which the informationally connected components are the single nodes, and after which the informationally connected component is the full graph. The emergence of a plateau $C=c_{0}>0$ for $r\rightarrow 2$ reflects the emergence of a progressively larger scale invariant structure. For $r\rightarrow 2$ we expect the plateau height to converge to the theoretical value $C\rightarrow c_{0}=1/2$ of the 1d-lattice passing by $c_{0}=4/6$ for random trees notoriously related to the spectral dimension $d=2c_{0}$ [3]. See also Section S4 illustrating the spectral entropy and specific heat curves for the three datasets independently depicted for each $r$-value.
@@
Remarkably, the thermodynamic efficiency is computed at the specific resolution scale $\tau\sim 10$ involved in the structural transition. Such critical resolution scale is identified by the bifurcation of the specific heat peaks depicted in panels (e)-(i) of Figure 4 and in panels (a)-(c) of Figure 5. In the seminal work [2], it is indicated to compute the thermodynamics efficiency, $\eta$, using as a reference the $\tau$-scale equal to the inverse of the Fiedler eigenvalue, i.e. the smallest non-zero Laplacian eigenvalue $\lambda_{2}=1/\tau_{diff}$, as this provides the time-scale where the diffusion process has reached equilibrium according to the authors, and then logarithmically divide $\tau_{diff}$ to progressively define the large propagation scales, the middle scales and the small scales.
@@
Nevertheless, $\tau_{diff}$ only approximates the end of the diffusion process which is better reflected by $\tau^{(2)}_{peak}$. Panels (a)-(c) of Figure 5 show the $\tau_{diff}$ in red next to the $\tau^{(1,2)}_{peaks}$ curves discussed above. The key difference is in the use of the ensemble average which employs all the eigenvalues to compute the specific heat. Thus, it is crucial to identify which is the $\tau$-scale which resolves the phase transition from the single trivial scale of the star graph to more complex topological traits of the amorphous network depicted in Figure 1. Such resolution scale is the relevant scale for the computation of the efficiency $\eta$ and defines an ”ultra-violet” cut-off in the characteristic frequency of the network across the transformation, similarly to what was discussed in the context of the Laplacian-Renormalization group [3]. In Section S3 the efficiency, $\eta$, is computed in the range interval indicated by the diffusion time as defined in [2] $\tau_{diff}=1/\lambda_{2}$, i.e. the red curves depicted in panels (a). Whether computed at those scales, $\eta$ shows a change in concavity and no consistent relation with the algorithmic accuracy.
@@
Moreover note in Figure 5 the slow down of diffusion indicated by the $\tau^{(2)}_{peak}$ curve as a function of $r$, we remark that $\tau^{(2)}_{peak}$ is a function of the parameter $r$ only which shapes the degree distribution as the number of nodes and edges is fixed across the transition. We remind that the Portegy’s algorithm allows us to stay in the regime of sparsity, symbolically $\delta(\frac{N-1}{|E|}-1)$.
@@
Figure 4: Thermodynamic & pattern-matching efficiency. In (a)-(c) the pattern-matching efficiency, $\theta$, is plotted as a function of $r$ for the three investigated datasets of increasing complexity (see arrow at the bottom of the figure). The shaded purple area indicates the tolerance $r$-range for the optimal pattern-matching efficiency, which includes points in one standard deviation from the global maximum indicated by the vertical red-dashed line. In (d)-(f) the optimal thermodynamic efficiency, $\eta$, as a function of $r$ overlaps with the pattern-matching efficiency. Vertical dashed-lines indicate the global maxima of the efficiencies. Remarkably, the thermodynamic efficiency is measured in the $\tau$-range of one order of magnitude from the bifurcation of the specific heat, $C$, depicted in panels (g)-(i). The $\tau$-resolution scale $\tau\sim 10$ identifies the relevant length scale resolving the transition from the single-peak to the emergence of a roughly plateau region, $C=c_{0}>0$, reflecting a roughly scale-invariant structure. Note that for datasets of higher complexity the highest efficiency, both thermodynamic and algorithmic, are posited at the critical value of the structural phase transition illustrated by the specific heat in panels (g)-(i). This is in line with the argument that states that a fairly simple dataset does not benefit from the critical point. Quantities are averaged over 50 independent simulations.
@@
Figure 5: Role of the ensemble average on the estimation of the $\tau$-scale. Panels (a)-(c) depict the $\tau^{(1,2)}_{peak}$ and $\tau_{diff}$ as a function of $r$. $\tau^{(1,2)}_{peak}$ are the $\tau$-location of the first and last peak of the specific heat in Figure 4. $\tau_{diff}$ is the Fiedler eigenvalue which provides a good estimation of the end of the diffusion process only for sufficiently large $r$. Note that the red curves retrace the profile of the upper branch of the specific heat for $\tau\gtrsim 10$ in panels (g)-(i) of Figure 4. At smaller $r$ values the equilibrium time is better reflected by $\tau^{(2)}_{peak}$. The equilibrium of the diffusion process is an increasing function of $r$. $\tau_{diff}$ curves are averaged over 50 independent simulations. The $\tau^{(1,2)}_{peak}$ curves are computed from the specific heat curves which are averaged over 50 independent simulations.
4.2 Performance at criticality

### v3

@@ 3 Methods
Our results are evaluated over subsets of four different datasets: the NIST [25], the MNIST [26], the FashionMNIST [27], and the CIFAR10 [28] datasets. These datasets contain randomly collected instances of image samples out of 10 classes. It should be noted that in this type of data, a pattern class can generally not be represented by a single centroid vector, but will contain multiple densities (’sub-style patterns’) such that the sample point cloud can be ordered naturally in a network structure.
Figure 2: Dataset complexity is estimated with multiple methods for each training dataset of 2000 samples encountered in this work. In green an estimate is given computing the Lempel-Ziv complexity [29], while in orange the intrinsic dimension is estimated with the Maximum Likelihood Estimation approach [30, 31, 32]. In blue the accuracy on the test set for a multilayer-perceptron (MLP) with two hidden layers of 128 units each. The left axis is referred to the quantities in green and orange, while the right axis is referred to the multilayer-perceptron accuracy. We observe that the CIFAR10 dataset is the most complex one, while the NIST dataset is the least complex among the four.
The MNIST database contains $24\times 24$ grey-scale images of handwritten digits [0-9] commonly used to benchmark machine- and deep-learning algorithms. The NIST dataset is the ancestor of the MNIST, it contains $16\times 16$ black-and-white images of the handwritten digits [0-9]. The FashionMNIST is a large dataset of $28\times 28$ gray-scale images of fashion products from 10 categories. The CIFAR-10 dataset contains $32\times 32$ color images belonging to 10 classes. In all our experiments a subset of 2000 samples is extracted from the train set and the test set independently. Samples from the FashionMNIST and MNIST are always binarized, while the samples from the CIFAR10 are normalized between 0 and 1 dividing each pixel value by 255, and values are rounded up to eighth decimal. In the Supplementary we show also the same results for larger networks generated from $4000$ samples from the MNIST and FashionMNIST. The 10 classes are equally represented both in the test and in the training set. The datasets used in our experiments exhibit different levels of complexity, see Figure 2. We estimate with multiple methods the dataset complexities on the same subset of 2000 training samples used in the rest of the manuscript. The first heuristic approach is based on the test accuracy of a small multilayer perceptron (2 layers with 128 units each). In pattern recognition, the complexity of a dataset is determined by the complexity of the curvilinear hyperplanes that would be needed to optimally separate patterns from different classes. A simple but effective parameter to gauge this facet of complexity is the accuracy that a classifier (such as the MLP) can achieve in determining the true class label for a data point. Thus, a more complex dataset would result in a lower accuracy for neural networks of fixed size when compared to a less complex dataset. The second method we use is the computation of the Lempel-Ziv complexity [29] of the aggregated training set. In this case, samples need to be binarized. The third method concerns the measure of the intrinsic dimension [15, 32] of the dataset using the Maximum Likelihood Estimation approach introduced in [30] with the correction proposed in [31]. A given dataset of gray-scale images of size $pixel\times pixel$ a priori lies on a high-dimensional space of $\mathbb{R}^{pix\times pix}$, nevertheless, because it might contain a specific class of objects that account for the "world" of our problem, for instance, handwritten digits, it spans a lower dimensional manifold. Even though the definition of such a manifold is an open problem, it is possible to estimate its dimension based on the distance of the $n$-neighbouring points [15, 32]. We set the hyper-parameter $n=3$.
While in Section 4 we discuss the full range of $r$, in Section 5 we limit the discussion to $r\leq 0.85$ as the accuracy, and hence the pattern-matching efficiency, quickly drops for larger $r$, moreover the algorithm shows instability as it doesn’t necessarily converge for $r\geq 0.9$. The Portegys algorithm we use is the extended version with $k=5$ introduced in [13] where the distance between patterns is the Euclidean distance.
Figure 3: Search method based on relative distance. In (a) the growth process is illustrated. At each step, a new pattern is linked to either the current parent pattern, $P$, or to one of its children, $C_{i}$. The parameter $r$ modulates the region of attraction of the child, indicated by the gray circle around $C_{1}$, and the one of the parent node. Points falling within the distance $r\cdot d(P,C_{i})$ are linked to the child, otherwise, an edge will be defined between the parent node $P$ and the new pattern. In (b) the search method based on relative distance is illustrated. Once a search pattern arrives at node $P$ it must choose to search either the subtree of $C_{1}$ or $C_{2}$. The choice is based on minimizing the distance to the expected patterns contained in the subtree of $C_{i}$. In the depicted case, patterns closer to the search pattern $S$ are expected to be in the subtree of $C_{2}$. Adapted from [33].
@@ 3.1 Pattern-matching efficiency
After the tree formation described in Section 1, the Portegys algorithm enters the search phase, where the best match to an unknown search pattern, $S$, is given using a first-best approach. The original goal of this algorithm was to restructure a list of examples in a tree-based way, such that the matching accuracy would not be much worse than nearest-neighbor matching on a full list while providing a considerable speed up due to the hierarchical organization into a tree-based structure. The growth process, described in Section 1 and depicted in panel (a) of Figure 3, ensures that subtrees of a given node have on average decreasing distances for decreasing hierarchical levels. This provides the possibility of using a search technique based on relative distances. The search method is depicted in panel (b) of Figure 3 and is an iterative process that begins at the root-node pattern. Once a search pattern, $S$, arrives at a certain parent node, $P$, it must decide to continue to search for the best match into one of its children subtree, $C^{\prime}_{1}$, $C^{\prime}_{2}$ in Figure 3. The choice is taken minimizing the Portegys search-distance: $d(S,C_{i})_{search}=d(S,C_{1})-r\cdot d(P,C_{i})=d(S,C^{\prime}_{i})$, which quantifies the distance between the $S$-pattern and the potential patterns appearing in the subtree of each child $C_{i}$, collectively indicated by the gray circle $C^{\prime}_{i}$ in panel (b). Where $d(\cdot)$ indicates any metric distance equipped with semi-positivity, symmetry, and triangular inequality properties. If $d(S,C_{i})_{search}$ is less than zero then it is set to zero. In panel (b) the bold segments in red indicate the search distance to each child node and equivalently the expected distance between the unknown pattern $S$ and their expected neighborhoods of patterns $C^{\prime}_{i}$.
It is natural to define a pattern-matching efficiency, $\theta$, in terms of class match accuracy, $A$, and the number of distances computed, $N_{dist}$:
@@
the logarithm is taken since the range of $N_{dist}$ can virtually span the interval $N_{dist}\in(0,N]$, where $N$ is the number of nodes. For the case of a star graph, $r\rightarrow 0$, the number of computed distances is maximum $N_{dist}=N$, for lower values of $r$ the number of computed distances decreases, potentially decreasing the class-match accuracy. When defining the pattern-matching efficiency we are making the simplified assumption that the pattern-matching algorithm runs on a physical substrate such that visiting any node has a fixed cost in terms of energy or terms of general resources. Hence, the efficiency $\theta$ quantifies the trade-off between the cost of used resources and the gain in performance, as depicted in panel (c, f) of Figure 4 for the MNIST and CIFAR10 respectively. Note in panel (d) that the CIFAR10 dataset has a peak for accuracy which we interpret by saying that for complex datasets (see also the case of the FashionMNIST in Section S7) some network structures do possess the ability to better encode information of the "world" in their traits. We postpone the discussion of this finding to Sections 5 and 6.
@@
Figure 4: Pattern-matching efficiency in (c, f) measured over the NIST and CIFAR10 respectively. The efficiency $\theta$ is defined as the trade-off between the classification accuracy, (a, d), and the logarithm of the number of distances computed to find the best match, (b, e). The dashed-red vertical line indicates the optimal $r$ value in terms of efficiency $\theta$. The shaded grey area designates the standard deviation. The purple shaded area in panel (c, f) indicates a tolerance $r$-domain of optimal efficiency. The tolerance domain is defined as the $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the maximum $\theta$, whose y-axis projection is indicated by the yellow shaded area. Curves are computed over 50 independent simulations with a subset of 2000 samples extracted from the NIST and CIFAR10 datasets. The 10-digit classes are equally represented both in the train- and test sets.
4 Phase portrait
Figure 5: Multiple phase transitions in hierarchical tree networks. In (a) the specific heat $C$ across the range $r\in[0,2]$. At $r\sim 1.35$, indicated by the dashed white line, the hierarchical network transits from the star graph to the amorphous configuration with the emergence of scale-invariant traits evidenced by a plateau in the specific heat. The scale-invariant region shows a second phase transition from amorphous to regular 1d-lattice at $r\simeq 1.05$ indicated by the dotted white line. For $r>0.35$, the network is always scale-invariant with decreasing spectral dimension up to the transition point, above which the spectral dimension saturates to $d_{s}=1=2\cdot c_{0}$, where $c_{0}$ is the plateau height in (c). Note that at the transition point $d_{s}=4/3$. In panel (c) the entropy (blue) and the specific heat (red) for $r=\{0.25,~{}0.35,~{}0.40,~{}0.70,~{}0.90,~{}1.05,2.00\}$. Dashed blue and green lines indicate the $\tau$-scale of the first and the second peak, $\tau^{(1,2)}_{peak}$, of the specific heat respectively. In (d), (e) the modularity and the normalized diameter, $\hat{d}=d/m$, together with their respective variances (dashed lines) identify the transition points. The modularity resolves the transition involving the mesoscale structure, while the diameter resolves the transition at the larger network scale. The legend indicates the two network sizes $N=500,~{}1000$. In (f) the difference of entropy at the two $\tau^{(1,2)}_{peak}$ peaks quantifies the normalized (log)number of informationally connected nodes across the scale-invariant $\tau$-scales, i.e. the number of nodes included in the scale-invariant region. Note the discontinuity at the transition from star to amorphous. The dotted line shows the peak of the variance at the transition point. In (g) the entropy surface is depicted and the entropy at the first and second specific heat peaks are illustrated, see legend. In (h) the $\tau$-scales of the two specific heat peaks are illustrated with blue and green dots respectively delimiting the three dynamical phases: the segregated phase (grey), the integrated phase (light yellow) and the green region where self-similar structures are integrated at a constant-rate. In (b) exemplary hierarchical trees of $N=20$ for growing $r$ from left to right are depicted illustrating the two phase transitions at the two resolution scales. All curves in this figure are realized for networks of $N=500$ randomly sampled from the FashionMNIST dataset and averaged over 50 independent realizations if not specified elsewhere.
The Portegys algorithm produces tree structures as those depicted in panel (b) of Figure 5 depending on the value of the control parameter $r$. Two structural phase transitions involving two different resolution scales can be identified (see panel (a)). The transition at the largest resolution scale can be described using as an order parameter a quantity that describes the large-scale network properties: $\hat{d}=d/m$, where $d$ is the diameter of the tree and $m$ is the number of edges equal to $m=N-1$. For hierarchical trees, another choice would be the normalized number of hierarchical levels $\hat{h}=h/(N-1)$ where $h$ is the number of hierarchical levels ranging from $h\in[0,N-1]$ (see illustration in panel (b)). Both quantities describe the "volume" expansion controlled by the parameter $r$. We chose the normalized diameter $\hat{d}$ and we identify this phase transition at $r\sim 1.05$ as indicated by the peak of the variance (dotted lines) in panel (e) for the two network sizes $N=500,~{}1000$ as indicated in the legend. The network structure transits from the amorphous regime to the regular 1d-lattice. Just below the transition point, a spectral dimension of $d_{s}=4/3=2c_{0}$ characteristic of random trees [34] is measured as indicated in (a) and shown in panels (c). Above the transition $d_{s}=1$, as it is expected for a 1d-lattice. Where $c_{0}$ is the specific heat plateau height.
The second change of phase is the transition from the star graph to the amorphous configuration at $r\sim 0.35$. Here, the emergence of scale-invariant structural properties is identified by the appearance of a second peak in the specific heat and a flat plateau in between the two $C$-peaks, as shown in panels (a) and (c) of Figure 5. Before the transition, a single peak in the specific heat identifies a single trivial $\tau$-resolution scale below which the informationally connected components are the single nodes, and after which the informationally connected component is the full graph. The emergence of a plateau $C=c_{0}>0$ for $r>0.35$ reflects the emergence of progressively larger scale-invariant structures that for $r\rightarrow 2$ would include the full network in the thermodynamic limit $N\rightarrow\infty$. This second change of phase is the one mainly investigated in this work as the peaks in the thermodynamic and pattern-matching efficiencies occur either in the vicinity of this transition border or are included in the scale-invariant regime. This transition, conversely to what we did with the amorphous to regular-lattice transition discussed above, can’t be identified by a quantity resolving the large-scale property of the network. In fact, we observe in panel (e) the lack of a peak in the variance of $\hat{d}$, which would indicate no sign of bona fide phase transition. Nevertheless, a phase transition indeed exists and an appropriate order parameter needs to be employed to identify it.
This phase change occurs at smaller resolution scales, specifically, the mesoscales involved in the structural transition from the star graph to the scale-invariant regime (see the bifurcation in the specific heat and the related quantities in panels (a,c,g,h)). Thus, we can identify this transition using as an order parameter the modularity [35] of the tree which we show in panel (e). We are taking advantage of the fact that the modularity for trees is bounded between $[0,1]$ and that in the case of a 1d-lattice, the modularity is equal to $1$ if we partition the network into $N$ communities, which is the case if we use the community detection algorithm introduced in [36]. We stress that modularity is a mesoscale descriptor while the diameter of the network is a large-scale descriptor. This motivates the choice of a mesoscale $\tau\sim 10$ employed in Section 5 that can resolve structural changes from the star graph phase to the amorphous scale-invariant phase.
In panel (f), the difference of entropy at the two $\tau^{(1,2)}_{peak}$ peaks counts the normalized (log)number of informationally connected nodes across the $\tau$-scales where the entropy is consumed at a constant rate, i.e. across the specific heat plateau. In simple terms, we count the fraction of nodes included in the scale-invariant region. We observe a discontinuity at the border of the scale-invariant transition that is not observed at the border of the amorphous to regular lattice transition.
As already discussed in Section 2.1, also along the $\tau$-direction a phase transition exists [22, 3] delimited by the two peaks of $C$ in $\tau^{(1,2)}_{peak}$ and, in between the two, we can distinguish three dynamical phases (see panel (h)). For $\tau<\tau^{(1)}_{peak}$, the network is in the informationally segregated phase; while for $\tau>\tau^{(2)}_{peak}$ is in the informationally integrated phase. Only for $r>0.35$ and $\tau^{(1)}_{peak}<\tau<\tau^{(2)}_{peak}$, we observe the dynamical regime of constant-rate entropy loss characteristic of scale-invariant structures with dilation symmetry [3, 23].
@@ 5 Results
Figure 6 shows the pattern-matching efficiency defined in Equation 6 and the thermodynamic efficiency of Equation 5 for the four investigated datasets.
@@
In panels (a)-(d) the pattern-matching efficiency, $\theta$, is depicted. The vertical red-dashed line indicates the $r$ value for the optimal trade-off, while the shaded purple region indicates a $r$-domain tolerance. The tolerance domain is defined as the $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the best trade-off.
In panels (e)-(h) the thermodynamic efficiency, $\eta$, is depicted and, for the case of the NIST and the FashionMNIST a notable consistency in the global maxima is observed with the pattern-matching efficiency $\theta$ in terms of optimal $r$. See the $r$-range spanned by the purple-shaded areas in panels (a)-(d) and the vertical dashed lines in panels (e)-(h). This consistency in the maxima of the two efficiencies suggests that thermodynamic or pattern-matching might be related to the context of this algorithm. In the Section 6 we discuss this matter.
@@
Similar to what would be the transition from vapor/liquid to an amorphous solid, at a critical value of $r$ topological complexity increases from the trivial traits of a star graph. Thus, the emergence of structure is reflected by the rise of a second peak in $C$ together with a roughly $C=c_{0}>0$ plateau in between the first and the last peak as is illustrated in panels (i)-(l) for the investigated dataset. Such transition is also depicted in Figure 7 by the bifurcation of the blue and green curves for $\tau^{(1)}_{peak}$ and $\tau^{(2)}_{peak}$ indicating the lower and upper branches in panels (i)-(l) of Figure 6. See also Section S4 illustrating the spectral entropy and specific heat curves for the four datasets independently depicted for each $r$-value.
@@
Remarkably, the thermodynamic efficiency is computed at the specific resolution scale $\tau\sim 10$ involved in the mesoscale structural transition from star graph to amorphous scale-invariant discussed in Section 4. Such critical resolution scale is identified by the bifurcation of the specific heat peaks depicted in panels (i)-(l) of Figure 6 and in panels (a)-(c) of Figure 7. In the seminal work [1], it is indicated to compute the thermodynamic efficiency, $\eta$, using as a reference the $\tau$-scale equal to the inverse of the Fiedler eigenvalue, i.e. the smallest non-zero Laplacian eigenvalue $\lambda_{2}=1/\tau_{diff}$, as this provides the time-scale where the diffusion process has reached equilibrium, and then logarithmically divide $\tau_{diff}$ to progressively define the large propagation scales, the middle scales and the small scales.
@@
Nevertheless, on the one hand, $\tau_{diff}$ only approximates the end of the diffusion process which is better reflected by $\tau^{(2)}_{peak}$. As shown in panel (h) of Figure 5 and in panels (a)-(d) of Figure 7 where the $\tau_{diff}$ in red next to the $\tau^{(1,2)}_{peaks}$ curves are illustrated. The key difference is in the use of the thermal ensemble average which employs all the eigenvalues to compute the specific heat. Thus, it is crucial to identify which is the $\tau$-scale that resolves the phase transition from the single trivial scale of the star graph to the more complex topological traits of the amorphous network depicted in panel (b) of Figure 5. Such resolution scale is the scale we consider of significant importance for the computation of the efficiency $\eta$ across the transition and defines an "ultra-violet" cut-off in the characteristic frequency of the network across the transformation, similarly to what was discussed in the context of the Laplacian-Renormalization group [3]. In Section S3 the efficiency, $\eta$, is computed in the range interval $\tau<10$ as is indicated by the diffusion time defined in [1] $\tau_{diff}=1/\lambda_{2}$, i.e. the red curves depicted in panels (a). Whether computed at those scales, $\eta$ shows a change in concavity and no relation with the pattern-matching accuracy.
@@
Moreover note in Figure 7 the slow down of diffusion due to the increasing "volume" of the network discussed in Section 4 and indicated by the $\tau^{(2)}_{peak}$ curve as a function of $r$, we remark that $\tau^{(2)}_{peak}$ is a function of the parameter $r$ only which shapes the degree distribution as the number of nodes and edges is fixed across the transition. We note that the Portegy’s algorithm allows us to stay in the regime of sparsity, symbolically $\delta(\frac{N-1}{|E|}-1)$.
@@
Figure 6: Thermodynamic & pattern-matching efficiency. In (a)-(d) the pattern-matching efficiency, $\theta$, is plotted as a function of $r$ for the four investigated datasets of increasing complexity (see arrow at the bottom of the figure). The shaded purple area indicates the tolerance $r$-range for the optimal pattern-matching efficiency, which includes points in one standard deviation from the global maximum indicated by the vertical red-dashed line. In (e)-(h) the optimal thermodynamic efficiency, $\eta$, as a function of $r$. In panels (f) and (h) a noticeable cohesion of the global maxima of the two efficiencies is observed. Vertical dashed lines indicate the global maxima of the efficiencies. Remarkably, the thermodynamic efficiency is measured in the $\tau$-range of one order of magnitude from the bifurcation of the specific heat, $C$, depicted in panels (i)-(l). The $\tau$-resolution scale $\tau\sim 10$ identifies the relevant length scale resolving the transition from the single-peak to the emergence of a roughly plateau region, $C=c_{0}>0$, reflecting a roughly scale-invariant structure. Note that for datasets of higher complexity the highest efficiency, both thermodynamic and pattern-matching, tends to be posited closer to the structural phase transition to scale-invariant illustrated by the specific heat in panels (i)-(l). This is in line with the argument that states that a fairly simple dataset does not benefit from the vicinity of the phase transition. Quantities are averaged over 50 independent simulations.
@@
Figure 7: Role of the ensemble average on the estimation of the $\tau$-scale. Panels (a)-(d) depict the $\tau^{(1,2)}_{peak}$ and $\tau_{diff}$ as a function of $r$. $\tau^{(1,2)}_{peak}$ are the $\tau$-location of the first and last peak of the specific heat in Figure 6. $\tau_{diff}$ is the inverse of the Fiedler eigenvalue which provides a good estimation of the end of the diffusion process only for sufficiently large $r$. Note that the red curves retrace the profile of the upper branch of the specific heat for $\tau\gtrsim 10$ in panels (i)-(l) of Figure 6. At smaller $r$ values the equilibrium time is better reflected by $\tau^{(2)}_{peak}$. The equilibrium of the diffusion process is an increasing function of $r$. $\tau_{diff}$ curves are averaged over 50 independent simulations. The $\tau^{(1,2)}_{peak}$ curves are computed from the specific heat curves which are averaged over 50 independent simulations.
5.1 Performance at the emergence of scale-invariance